OpenStudy (anonymous):

Evaluate (3x)^2 - 3x^2 when x = 5.

OpenStudy (idealist10):

(3*5)^2-3(5)^2=15^2-3*25=225-75=150. Answer: 150

OpenStudy (anonymous):

ok im only in the 7th grade and i have never see '^' befor so im i know i cant help u sorry id try but i dont want u to get it wrong

OpenStudy (anonymous):

ok well that's fine

OpenStudy (anonymous):

helpmeplzzzzz can you help me with Simplify 2y + 3y+4y

OpenStudy (idealist10):

^ means exponent. for example, 5^2=5*5=25. 2y+3y+4y=5y+4y=9y Answer: 9y You just combine like terms.

OpenStudy (anonymous):

thanks@Idealist10